# Unifying Search and Recommendation: A Generative Paradigm Inspired by Information Theory

### v2

@@ 5.3. Empirical analysis
In our experiments, we calculate mutual information for both S&R tasks under the discriminative and generative paradigms. For each task, we estimate $I(X;Y_{t})$ by sampling data from the joint distribution $P(X,Y_{t})$ and comparing it with samples from the independent marginal distributions $P(X)$ and $P(Y_{t})$. This process provides a quantitative measure of how well the input and output distributions are aligned in each paradigm. We compare GenSR with the most powerful discriminative paradigm baseline UniSAR in two public datasets: Amazon and KuaiSAR (see detailed dataset description and baseline description in Section 6.1). The results, presented in Table 1, show that the generative paradigm consistently achieves higher mutual information for both tasks. This improvement aligns with our theoretical analysis, which is attributed to the task-specific subspaces.
6. Experiments
In this section, we conduct a comprehensive experimental study to address the following research questions. RQ1: How does the performance of GenSR compare to other baseline methods across various datasets in both recommendation and search scenarios? RQ2: How well can GenSR enhance mutual information compared to discriminative approaches? RQ3: What is the impact of different strategies (e.g., dual-view representation and S&R history soft filtering) within GenSR on its overall performance?
Table 2. Overall performance of GenSR and other baselines under recommendation scenario. The best and second-best results are highlighted in bold and underlined fonts, respectively. * denotes statistically significant improvements (t-test, $p$ 0.01).
Recommendation Joint Search and Recommendation
Dataset Metric DIN GRU4Rec SASRec BERT4Rec FMLP CoLLM IDGenRec NRHUB SeqRec SESRec JSR USER UnifiedSSR UniSAR BSR GenSR
Amazon R@5 0.5170 0.4949 0.5295 0.5311 0.5356 0.5800 0.2305 0.4988 0.5401 0.5623 0.5467 0.5441 0.5196 0.5874 0.1417 0.6102*
R@10 0.6525 0.6548 0.6772 0.6658 0.6879 0.6811 0.2818 0.6503 0.6860 0.6864 0.6779 0.6854 0.6707 0.7020 0.2278 0.7200*
N@5 0.3726 0.3388 0.3747 0.3954 0.3739 0.4651 0.1791 0.3487 0.3859 0.4245 0.3970 0.3964 0.3662 0.4513 0.0899 0.4922*
N@10 0.4165 0.3907 0.4225 0.4390 0.4232 0.5001 0.1955 0.3977 0.4333 0.4648 0.4396 0.4422 0.4151 0.4885 0.1177 0.5278*
KuaiSAR R@5 0.4509 0.3764 0.4065 0.3699 0.4292 0.4092 0.3862 0.3920 0.4956 0.4791 0.4086 0.3981 0.5169 0.2180 0.5235*
R@10 0.6179 0.5788 0.6007 0.5885 0.6159 0.5915 0.5610 0.5890 0.6643 0.6453 0.5627 0.5939 0.6792 0.3395 0.6819*
N@5 0.3104 0.2345 0.2671 0.2381 0.2851 0.3199 0.2572 0.2552 0.3432 0.3315 0.2820 0.2617 0.3632 0.1472 0.3686*
N@10 0.3643 0.3087 0.3298 0.3083 0.3453 0.3683 0.3136 0.3186 0.3978 0.3853 0.3318 0.3249 0.4158 0.1860 0.4199*
Table 3. Overall performance of GenSR and other baselines under search scenario. Bold indicates the best result and underline the second-best, with * denoting statistically significant improvements (t-test, $p$ 0.01).
Search Joint Search and Recommendation
Dataset Metric QEM HEM AEM ZAM TEM CoPPS JSR USER UnifiedSSR UniSAR BSR GenSR
Amazon R@5 0.7100 0.6778 0.7095 0.7109 0.8185 0.8169 0.7038 0.7631 0.7744 0.8190 0.4786 0.9782*
R@10 0.8186 0.8267 0.8443 0.8468 0.9051 0.9051 0.8225 0.8697 0.8812 0.8977 0.6015 0.9664*
N@5 0.5066 0.4736 0.5114 0.5147 0.6303 0.6281 0.5173 0.6000 0.5847 0.6875 0.3485 0.9853*
N@10 0.5422 0.5221 0.5554 0.5590 0.6587 0.6570 0.5563 0.6348 0.6196 0.7132 0.3883 0.9687*
KuaiSAR R@5 0.6020 0.6505 0.5956 0.6117 0.6502 0.6616 0.7162 0.7304 0.7377 0.7476 0.1573 0.7661*
R@10 0.7182 0.7653 0.7182 0.7344 0.7632 0.7707 0.7961 0.8149 0.8320 0.8369 0.2075 0.8372
N@5 0.4575 0.5029 0.4415 0.4560 0.4887 0.4977 0.5962 0.6069 0.5991 0.6417 0.1359 0.7564*
N@10 0.4953 0.5400 0.4812 0.4959 0.5254 0.5331 0.6221 0.6342 0.6297 0.6708 0.1520 0.7655*
6.1. Experimental setups
6.1.1. Datasets.
We evaluate GenSR on the two publicly accessible datasets that contain both S&R histories. 1) Amazon [^3] https://jmcauley.ucsd.edu/data/amazon/. is a recommendation review dataset from Amazon. We use the ”Kindle Store” 5-core sub-dataset from the dataset, which contains user reviews and ratings for Kindle books. Adopting the procedure from prior work (Si et al., 2023; Shi et al., 2024), we first generate synthetic search behaviors and queries based on this recommendation dataset and then apply a leave-one-out strategy for dataset splitting. The detailed process for generating search behaviors is described in (Ai et al., 2017). This automatically constructed search dataset has been widely adopted by the product search community and prior work on unifying search and recommendation (Si et al., 2023; Shi et al., 2024; Ai et al., 2017; Van Gysel et al., 2016; Guo et al., 2019). It contains 68,223 users, 61,934 items, 4,298 queries, 989,618 recommendation interactions, and 934,664 search interactions. 2) KuaiSAR [^4] https://kuaisar.github.io/. is a micro-video dataset that offers authentic S&R behaviors of users, featuring detailed item categories and real user queries. We preprocess the data to include only users showing both S&R behaviors. Following (Shi et al., 2024), we filter for users and items with at least five interactions and allocate the latest day’s data for testing, the previous day for validation, and the earlier data for training. It contains 25,877 users, 6,890,707 items, 453,667 queries, 14,605,716 recommendation interactions, and 5,059,169 search interactions.
6.1.2. Baselines.
The recommendation baselines are: 1) DIN (Zhou et al., 2018) generate ad-specific user interest representations from historical behaviors. 2) GRU4Rec (Hidasi et al., 2016) utilizes RNNs for session-based recommendations. 3) SASRec (Kang and McAuley, 2018) applies self-attention to model long-term user behavior. 4) BERT4Rec (Sun et al., 2019) uses bidirectional modeling to predict masked items in sequences. 5) FMLP (Zhou et al., 2022) uses an all-MLP model that applies frequency-domain filtering to sequence recommendation. 6) CoLLM (Zhang et al., 2023) integrates collaborative embeddings into LLMs for recommendation. 7) IDGenRec (Tan et al., 2024) generates personalized recommendations using textual item IDs. [^5] We only implement IDGenRec on Amazon due to its requirement for explicit semantic textual information of items, which is unavailable for KuaiSAR.
The search baselines are: 8) QEM (Ai et al., 2019) assesses item-query matching scores. 9) HEM (Ai et al., 2017) utilizes a hierarchical embedding model that captures the semantics of queries, items, and users. 10) AEM (Ai et al., 2019) offers personalizes search results through an attention mechanism based on query and user data. 11) ZAM (Ai et al., 2019) enhances AEM by appending a zero vector for better item differentiation. 12) TEM (Bi et al., 2020) implements a Transformer-based embedding model for personalized product search. 13) CoPPS (Dai et al., 2023) introduces contrastive learning to refine search outcomes from sparse and noisy data.
The unifying baselines include two categories: (a) Search enhanced recommendation: 14) NRHUB (Wu et al., 2019) enhances user models with search queries for improved recommendation. 15) SeqRec (He et al., 2022) uses queries to drive user behavior modeling with transformers. 16) SESRec (Si et al., 2023) uses contrastive learning to sort user interests from S&R data. (b) Jointly unified S&R: 17) JSR (Zamani and Croft, 2018) jointly models S&R tasks with a shared neural network. 18) USER (Yao et al., 2021) merges user behaviors from both domains into a cohesive sequence for joint modeling. 19) UnifiedSSR (Xie et al., 2024) employs an intent-oriented session modeling module to infer user intent. 20) UniSAR (Shi et al., 2024) models user behavior transitions between S&R. 21) BSR (Penha et al., 2024) employs atomic item IDs to do the S&R task in generative model.
6.1.3. Experimental settings.
Following previous work (Shi et al., 2024; Xie et al., 2024; Yao et al., 2021), we randomly sample 99 negative items that the user has not interacted with and combine them with the ground truth item to create the candidate item list. All the models are then tasked with ranking these candidate items to evaluate top-K S&R performance using Recall@K and NDCG@K with $K=\{5,10\}$. To ensure a fair comparison, we implement constrained beam search during inference for two full-ranking generative baselines, IDGenRec and GenSR, which guarantees that the generated results are restricted to the valid candidate set. Regarding our proposed GenSR, we select flan-t5-base as the backbone. The best hyper-parameters are selected with the searching scopes as follows: the temperature factor $\tau$ and the loss balance factor $\beta$ are tuned within $\{0.01,0.03,0.05,0.07,0.09\}$ and $\{0.001,0.01,0.1,0.15,0.2,0.25,0.3\}$.
6.2. Overall performance (RQ1)
We conduct experiments in both S&R scenarios. The results, illustrated in Table 2 and Table 3, yield several key observations:
  • GenSR consistently outperforms all baselines across all datasets in both S&R tasks. This superior performance can be attributed to GenSR’s dual representation learning and S&R task unifying under the superior generative paradigm.
  • Unified methods for S&R generally outperform baselines that focus on only one task. This highlights that leveraging shared user behaviors and item interactions in the S&R scenario can mutually enhance performance in both tasks.
  • The performance of IDGenRec is relatively weaker, potentially due to its reliance on semantic information for the recommendation task, which may limit its ability to capture collaborative filtering information explicitly. Similarly, BSR’s performance may be constrained by its use of atomic item IDs, which lack semantic information to enrich item representations further.
6.3. In-depth analysis
@@
6.3.1. Parameter space visualization analysis (RQ2)
We analyze the parameter space to support three insights: (1) S&R tasks exhibit distinct data distributions; (2) discriminative paradigm forces both tasks into a shared space, causing gradient conflicts; and (3) generative paradigm naturally separates this space, reducing such conflicts. Experiments are based on KuaiSAR, the only real-world public dataset. To examine the raw data distributions of S&R tasks, we utilize Kernel Density Estimation (KDE) to approximate $P(X_{S},Y_{S})$ and P$(X_{R},Y_{R})$ (Terrell and Scott, 1992), using concatenated $X$ and $Y$ vectors. After dimensionality reduction via t-SNE(Van der Maaten and Hinton, 2008), we compute KDE in the embedded space. The resulting contours (Figure 3a) reveal distinct patterns, confirming distributional differences of two tasks. Given the observed differences in task distributions, we then examine parameter space overlap under the discriminative setting by projecting task-specific hidden states using PCA (Figure 3b). The significant overlap reflects optimization conflicts due to mismatched task distributions. In contrast, the generative paradigm naturally partitions the parameter space into task-specific subspaces that align with the separate distributions of S&R tasks (Figure 3c). This separation enables independent optimization for each task, effectively mitigating task conflicts and enhancing the model’s ability to handle both tasks within a unified framework.
@@ 6.3.1. Parameter space visualization analysis (RQ2)
6.3.2. Ablation study (RQ3)
We conduct an ablation study to verify the effectiveness of each part of GenSR. We assess the unique contributions of dual-view representation, S&R history soft filtering, and contrastive learning. The results, presented in Figure 4, reveal the following findings: 1) Preserving only one representation view reduces performance on both datasets. This result underscores the advantage of independently learning item representations from the CF and semantic views, enabling the model to effectively integrate the diverse information needed to support both S&R tasks. 2) Excluding S&R soft filtering also leads to a decline in performance, which demonstrates the effectiveness of prioritizing relevant information from interaction histories. 3) Removing contrastive learning similarly decreases performance, which highlights the importance of aligning the CF-view and semantic view for the same user.
Figure 5. Hyper-parameters analysis under Amazon.
6.3.3. Hyper-parameter analysis (RQ3)
We examine the impact of two hyper-parameters of GenSR in Figure 5 using the Amazon dataset (excluding KuaiSAR due to similar trends): the temperature factor $\tau$ used in contrastive learning and the loss balance factor $\beta$ controlling the weight of the contrastive loss relative to the S&R task loss. The result shows that: 1) Varying $\tau$ affects how sharply similarities are emphasized among CF and semantic embeddings. Smaller $\tau$ values highlight high-similarity pairs more strongly but risk overfitting, while larger $\tau$ values foster more even attention at the expense of less pronounced discrimination. 2) Similarly, $\beta$ determines the trade-off between focusing on embedding alignment versus optimizing the main S&R tasks. An excessively high $\beta$ could overemphasize embedding alignment, marginalizing S&R performance; a too-low $\beta$ may fail to leverage the complementary strengths of the two views.
@@ 7. Conclusion
In this work, we have introduced GenSR, a novel generative paradigm designed to unify search and recommendation tasks, which leverages task-specific prompts to partition the model’s parameter space into subspaces, thereby enhancing mutual information. To construct effective subspaces for each task, GenSR first prepares informative representations for each subspace through dual representation learning and then optimizes both subspaces in one unified model. Comprehensive experiments on two public datasets have validated GenSR’s effectiveness, showing consistent improvements over state-of-the-art methods across S&R tasks.
@@
Appendix
Proof of Lemma 5.1
We consider two paradigms: the discriminative paradigm with parameter $w_{\theta}$ and the generative paradigm with parameter $w_{\phi}$. Let the true parameters be $w_{\theta}^{*}$ and $w_{\phi}^{*}$, and let their estimates from data be $\hat{w}_{\theta}$ and $\hat{w}_{\phi}$ respectively.
The parameter approximation is obtained by minimizing their respective loss functions: [^6] For binary generation tasks (e.g., “yes” vs. “no”), the negative log-likelihood in the generative setting aligns with the binary cross-entropy loss of the discriminative setting, allowing the same format of loss.
@@
In the binary classification setting, it has been proved that the discriminative parameter can be represented as the difference of the generative parameters across classes (i.e., $w_{\theta}^{*}=w_{\phi,1}^{*}-w_{\phi,0}^{*}$, where $w_{\phi,0}^{*}$ and $w_{\phi,1}^{*}$ denote the true parameters for class 0 and class 1, ) (Prasad et al., 2017). Hence, comparing parameter approximation errors for the two paradigms is meaningful. To do that, we expand the loss functions around the respective true parameters using Taylor’s expansion. We define the approximation errors as follows: $\Delta_{D}=\hat{w}_{\theta}-w_{\theta}^{*}$, $\Delta_{G}=\hat{w}_{\phi}-w_{\phi}^{*}$. For each paradigm, we have:
@@
where $\gamma_{G}$ and $\gamma_{D}$ are the separability constants for the generative and discriminative paradigms respectively. Since it has been proven that the generative paradigm enjoys higher separability (i.e. $\gamma_{G}>\gamma_{D}$(Prasad et al., 2017), we have $\|R_{G}(\Delta)\|_{1}<\|R_{D}(\Delta)\|_{1}$, for any fixed $\Delta$.
Assuming that the perturbations are small enough for the linear part of the expansion to dominate, we can apply the Taylor expansions:
@@
Moreover, since the generative paradigm benefits from both explicit partitioning by prompts and higher separability, its Hessian takes a block-diagonal form and is better conditioned: [^7] Although this uniform bound is stated for an arbitrary (fixed) $\Delta$, in practice the error analysis is applied to the respective perturbations $\Delta_{D}$ and $\Delta_{G}$ in their own local neighborhood. Since both $\Delta_{D}$ and $\Delta_{G}$ are assumed to be small (and we are only using the bound locally), it is reasonable to apply the same type of bound separately to each—even though they are different. The uniformity of the bound assures us that the same structure holds regardless of which (small) perturbation is under consideration.
@@
Consider a Gaussian model given by: $Y=w^{*}X+b^{*}+\epsilon$, with $\epsilon\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$ independent of $X$, $X\sim\mathcal{N}(\mu_{X},\sigma_{X}^{2})$. Building on prior work in information theory and statistical learning (e.g.,(Leung and Barron, 2006; Gabrié et al., 2018)), mutual information between input $X$ and prediction $\hat{Y}$, parameterized by model parameter $w$, is expressed as:
(14) $I(X;\hat{Y})=\frac{1}{2}\ln\left(1+\frac{w\sigma_{X}^{2}}{\sigma_{\epsilon}^{2}}\right).$
Suppose the noise variance $\sigma_{\epsilon}$ is the same in both paradigms (Hastie et al., 2009), as parameter estimates $\hat{w}$ approach the true parameter $w^{*}$, the mutual information $I(X;\hat{Y})$ approaches the theoretical upper bound $I(X;Y)$, which is identical for both paradigms. Given that we have established Eq.(13) and Eq.(14), the generative paradigm achieves lower approximation error, leading directly to higher mutual information:

### v3

@@ 6.4.1. Quantifying mutual information gains
In our experiments, we calculate mutual information for both S&R tasks under the discriminative and generative paradigms. For each task, we estimate $I(X;Y_{t})$ by sampling data from the joint distribution $P(X,Y_{t})$ and comparing it with samples from the independent marginal distributions $P(X)$ and $P(Y_{t})$. This process provides a quantitative measure of how well the input and output distributions are aligned in each paradigm. We compare GenSR with the most powerful discriminative paradigm baseline UniSAR in two public datasets: Amazon and KuaiSAR (see detailed dataset description and baseline description in Section 6.1). The results, presented in Table 5, show that the generative paradigm consistently achieves higher mutual information for both tasks. This improvement aligns with our theoretical analysis, which is attributed to the task-specific subspaces. We next provide a parameter-space visualization to complement these quantitative results.
6.4.2. Parameter space visualization analysis
We empirically analyze the model’s parameter space to deeply investigate and visually support our core claim: the proposed generative paradigm can effectively enhance the mutual information between the model input $X$ and task-specific output $Y_{t}$. We present our analysis in three steps: (1) the inherent differences in the original data distributions of the search and recommendation tasks; (2) the optimization challenges posed by the discriminative paradigm under a shared parameter space, which may lead to low mutual information; and (3) how the generative paradigm naturally achieves high mutual information by partitioning the parameter space. All analyses are conducted based on the KuaiSAR dataset, as it is the only real-world, publicly available dataset, offering authentic complexity and user behavior patterns and avoiding biases that may be introduced by synthetic data, thereby making our observations more representative.
[…]
@@ 6.4.2. Parameter space visualization analysis
Given the observed differences in task distributions, we further investigate the parameter space when unifying S&R tasks under the discriminative paradigm. Specifically, we extract the model’s hidden states for S&R tasks and use Principal Component Analysis (PCA) to visualize the activation distribution in two dimensions as shown in Figure 3(b). The results indicate that the hidden state distributions of S&R tasks under the discriminative paradigm exhibit significant overlap. This suggests that the model is forced to learn and represent two fundamentally different tasks within a single shared parameter space. Such enforced sharing leads to limitations in representational capacity: the model struggles to learn sufficiently distinct, task-specific representations for each task. From an information-theoretic perspective, this overlap causes the task-relevant information in the input $X$ to become entangled, thereby limiting the upper bound of mutual information. When the model cannot effectively differentiate or capture the distinct patterns of each task, it becomes difficult to find optimal gradient directions during training, which is the root cause of gradient conflicts (which will be quantified in Section 6.5).
Finally, we analyze the parameter space under our generative paradigm in Figure 3(c). The generative paradigm naturally partitions the parameter space into task-specific subspaces that align with the separate distributions of S&R tasks. This separation yields multiple advantages: (1) Enhanced Mutual Information: The independence of subspaces ensures that input features are mapped more clearly and accurately to their corresponding task outputs. This allows each task’s representation learning and optimization to proceed without interference from the other, significantly enhancing the mutual information between input $X$ and task-specific output $Y_{t}$. This enriched mutual information translates directly into stronger task-specific representation capacity and improved prediction accuracy. (2) Mitigated Gradient Conflicts: By partitioning the space, the optimization gradients for the search and recommendation tasks can evolve independently within their own ”dedicated” subspaces. This fundamentally reduces direct conflicts between task-specific optimization goals, enabling the model to converge more stably and efficiently while maximizing performance for both tasks.
6.5. Gradient conflict analysis (RQ3)
In this section, we further empirically evaluate the degree of gradient conflict in both the discriminative paradigm and our proposed method to support that our method alleviates gradient conflict compared to the discriminative paradigm.
To assess this effect, we compare our approach with the strongest discriminative baseline, UniSAR. At each training step, both models are supplied with the same mini-batch. We compute the loss for the search task and the recommendation task separately for the current mini-batch. Then, from both models (GenSR and UniSAR), we extract the corresponding gradients from the same self-attention layer in the transformer encoder. This layer is typically crucial for learning high-level representations and is where task-shared information is most concentrated, making it a likely site for gradient conflict. And then, the extracted high-dimensional gradient tensors are flattened into one-dimensional vectors. We compute the cosine similarity between the two flattened gradient vectors. A higher cosine value indicates less conflict between the two tasks.
Figure 4 shows the similarity distribution across the training process. The key trends and conclusions are as follows:
  1. (1)
    Fluctuations and Negative Similarity in UniSAR: As a representative of the discriminative paradigm, UniSAR exhibits dramatic fluctuations in its gradient similarity curve, frequently dipping into the negative range. These sharp fluctuations indicate frequent disagreements between the optimization objectives of S&R tasks over shared parameters across training steps (mini-batches). Additionally, the recurring dips into negative values (often around or below -0.5) clearly point to severe gradient conflicts. This suggests that at these steps, the gradient directions for search and recommendation are adversarial. Such opposing updates cause the model parameters to oscillate between conflicting objectives, slowing convergence, reducing training efficiency, and ultimately resulting in suboptimal performance.
  2. (2)
    Sustained Positive and Smooth Curve in GenSR: In contrast, GenSR maintains a consistently positive has a smaller amplitude in the gradient similarity curve throughout training. The predominantly positive cosine similarities indicate that GenSR rarely encounters task conflicts; in most cases, the gradients of search and recommendation tasks are aligned or at least non-interfering. This means that even when the two tasks pursue different objectives, their updates to shared parameters remain largely compatible. Additionally, the smoother evolution of the curve further reflects training stability and efficiency. In the absence of severe gradient opposition, model parameters converge more steadily toward globally optimal solutions.
These contrasting trends clearly demonstrate that assigning search and recommendation to distinct subspaces significantly reduces gradient conflict and enables more stable optimization compared to the discriminative paradigm.
Figure 4. Gradient Conflict Analysis.
@@ 6.5. Gradient conflict analysis (RQ3)
6.6. Ablation study (RQ4)
To systematically assess the individual contributions of each critical module within GenSR, we conducte a comprehensive ablation study. Specifically, our aim is to quantitatively analyze the independent impact of the dual-view representation learning, S&R history soft filtering, and contrastive learning modules on the model’s overall performance. Figure 5 presents the ablation experimental results, revealing the following key findings:
  1. (1)
    Contribution of Dual Representation Learning: When the model is restricted to a single representation view (i.e., utilizing only the CF view or the semantic view), there is a substantial performance degradation across both recommendation and search tasks on both datasets. For instance, on the Amazon dataset, the NDCG@5 for the recommendation task dropped by approximately 26.5% compared to the full model, while for the search task, NDCG@5 decreased by about 6.1%. The KuaiSAR dataset exhibited similar performance drops. These findings underscore the intrinsic complementarity between CF and semantic information for unifying S&R tasks. By independently learning representations from these two distinct views, GenSR effectively integrates the diverse information needed to support both S&R tasks.
  2. (2)
    Contribution of S&R History Soft Filtering: Excluding the S&R history soft filtering module led to a noticeable decline in model performance. On the Amazon dataset, the NDCG@5 for the recommendation and search tasks decreased by approximately 12.2% and 3.1%, respectively. These results underscore the critical role of this filtering strategy in extracting relevant information from complex interaction histories.
  3. (3)
    Contribution of Contrastive Learning: The removal of the contrastive learning component resulted in a significant decline in model performance. On the Amazon dataset, the NDCG@5 for the recommendation and search tasks dropped by approximately 6.1% and 1.0%, respectively. These findings highlight the critical role of contrastive learning in aligning representations from different views. While dual-view representation learning independently captures CF and semantic information, contrastive learning acts as a crucial mechanism to integrate these complementary representations within a shared embedding space. By encouraging the CF and semantic embeddings of the same user to be spatially closer while separating embeddings of different users, contrastive learning ensures both views jointly contribute to robust and comprehensive user preference modeling. This cross-view alignment is essential for enhancing the model’s generalization capability, as it prevents potential representation misalignment or disentanglement that may arise from independent learning. Ultimately, this mechanism maximizes the synergistic benefits of leveraging information from dual sources.
Figure 6. Hyper-parameter analysis for the Amazon dataset.
6.7. Hyper-parameter analysis (RQ4)
To thoroughly understand the robustness and optimal configuration of GenSR, we conducte a sensitivity analysis on two critical hyperparameters: the temperature factor ($\tau$) within the contrastive learning objective, and the loss balance factor ($\beta$) that governs the weighting of the contrastive loss relative to the main S&R task loss. This analysis is performed on the Amazon dataset, with observed trends being consistent across the KuaiSAR dataset, thus the Amazon results are presented for conciseness in Figure 6. Our findings are as follows:
  1. (1)
    The temperature factor $\tau$ is a critical parameter in shaping the similarity distribution within the contrastive learning objective. Specifically, it scales the logits before the softmax operation, thereby controlling the sharpness of the similarity distribution. A smaller $\tau$ leads to a sharper distribution, assigning higher confidence to truly similar pairs and emphasizing fine-grained distinctions between positive and negative samples. Conversely, a larger $\tau$ results in a smoother distribution, which softens the penalty for less similar negative pairs and can reduce the discriminative power. As shown in the Figure 6(a), GenSR’s performance, measured by NDCG@5 for both recommendation and search tasks, exhibits a clear peak within a specific range of $\tau$. When $\tau$ is too small (e.g., 0.01), performance significantly degrades (e.g., NDCG@5 for recommendation drops by approximately 1.5% from peak). This is likely due to the contrastive loss becoming overly sensitive to minor differences or noise, potentially leading to feature collapse or focusing too intensely on easy negative samples, hindering the learning of generalized representations. Conversely, excessively large $\tau$ values (e.g., 0.09) also lead to performance drops (e.g., NDCG@5 for recommendation drops by approximately 0.8% from peak), as the loss function becomes less discriminative, failing to adequately distinguish between positive and negative pairs. This indicates that finding a balanced $\tau$ is critical for effective contrastive learning. An appropriate $\tau$ ensures that the model can sufficiently distinguish between positive (CF and semantic views of the same user) and negative samples (views of different users) without becoming overly sensitive or under-discriminating.
  2. (2)
    The loss balance factor determines the relative importance of the contrastive learning loss $L_{c}$ compared to the primary S&R task generation loss $L_{S\&R}$. It controls the trade-off between aligning the dual-view representations and optimizing for accurate task-specific outputs. Figure 6(b) illustrates the impact of $\beta$ on performance. We observe that performance for both S&R tasks generally improves as $\beta$ increases from a very low value (e.g., 0.001) and then stabilizes or slightly declines after reaching an optimal range. For instance, when $\beta$ is too low (e.g., 0.001), the NDCG@5 for recommendation is significantly lower, suggesting that the contrastive loss has insufficient influence to properly align the CF and semantic views, thus failing to leverage their complementary strengths. As $\beta$ increases into the optimal range, the model benefits from the improved aligned representations. However, excessively high $\beta$ values can shift the model’s focus disproportionately toward embedding alignment, potentially marginalizing the primary S&R task objectives. This analysis underscores the importance of carefully selecting $\beta$. A well-chosen value ensures the contrastive loss contributes sufficiently to robust representation learning and cross-view alignment without overshadowing the core S&R objectives, maintaining a balance that supports both embedding alignment and task-specific accuracy.
@@ 7. Conclusion
In this work, we have introduced GenSR, a novel generative paradigm that unifies search and recommendation by leveraging task-specific prompts to partition the model’s parameter space into specialized subspaces. This subspace partitioning framework encourages each task to learn distinct yet complementary representations, thereby enhancing mutual information and reducing gradient conflict within a unified model. To enable effective subspace modeling, GenSR first constructs dual-view user representations that capture both collaborative and semantic signals across S&R behaviors. These signals are aligned through contrastive learning and are fused with prompt-guided generation, allowing the model to adaptively generate task-specific outputs based on a unified interaction history. Unlike previous discriminative methods that rely heavily on manual architecture design or rigid parameter sharing, GenSR offers a more flexible and expressive alternative grounded in information-theoretic principles.
Extensive experiments on two public datasets, Amazon and KuaiSAR, demonstrate the effectiveness of our approach. GenSR consistently outperforms strong discriminative and generative baselines on both search and recommendation tasks. In addition to superior ranking performance, our analysis shows that GenSR achieves higher estimated mutual information and better gradient alignment, confirming that the proposed method improves the model representation capability and mitigate gradient conflict between two tasks.
@@
Appendix (Proof of Theorem 5.1)
This appendix provides a detailed proof of Theorem 4.1, aiming to theoretically demonstrate how our proposed generative paradigm, GenSR, effectively enhances mutual information in unifying S&R tasks.
We consider two paradigms for unifying S&R: the discriminative paradigm with parameter $w_{\theta}$ and the generative paradigm with parameter $w_{\phi}$. Let the true parameters be $w_{\theta}^{*}$ and $w_{\phi}^{*}$, and let their estimates from data be $\hat{w}_{\theta}$ and $\hat{w}_{\phi}$ respectively. The parameter approximation is obtained by minimizing their respective loss functions: [^7] For binary generation tasks (e.g., “yes” vs. “no”), the negative log-likelihood in the generative setting aligns with the binary cross-entropy loss of the discriminative setting, allowing the same format of loss.
@@
In the binary classification setting, it has been proved that the discriminative parameter can be represented as the difference of the generative parameters across classes (i.e., $w_{\theta}^{*}=w_{\phi,1}^{*}-w_{\phi,0}^{*}$, where $w_{\phi,0}^{*}$ and $w_{\phi,1}^{*}$ denote the true parameters for class 0 (“yes”) and class 1 (“no”) ) (Prasad et al., 2017). Hence, comparing parameter approximation errors for the two paradigms is meaningful. To do that, we expand the loss functions around the respective true parameters using Taylor’s expansion. We define the approximation errors as follows: $\Delta_{D}=\hat{w}_{\theta}-w_{\theta}^{*}$, $\Delta_{G}=\hat{w}_{\phi}-w_{\phi}^{*}$. For each paradigm, we have:
@@
where $\gamma_{G}$ and $\gamma_{D}$ are the separability constants for the generative and discriminative paradigms respectively. Since it has been proven that the generative paradigm enjoys higher separability (i.e. $\gamma_{G}>\gamma_{D}$) (Prasad et al., 2017), we have $\|R_{G}(\Delta)\|_{1}<\|R_{D}(\Delta)\|_{1}$, for any fixed $\Delta$.
Assuming that the perturbations are small enough for the linear part of the expansion to dominate, we can apply the Taylor expansions (Song et al., 2011):
@@
Furthermore, the generative paradigm, through its explicit partitioning of the parameter space via task-specific prompts, and its inherent higher separability, results in a Hessian matrix with better conditioning during optimization: [^8] Although this uniform bound is stated for an arbitrary (fixed) $\Delta$, in practice the error analysis is applied to the respective perturbations $\Delta_{D}$ and $\Delta_{G}$ in their own local neighborhood. Since both $\Delta_{D}$ and $\Delta_{G}$ are assumed to be small (and we are only using the bound locally), it is reasonable to apply the same type of bound separately to each—even though they are different. The uniformity of the bound assures us that the same structure holds regardless of which (small) perturbation is under consideration.
@@
This indicates that the generative paradigm can achieve a lower parameter approximation error, meaning its estimated parameters are closer to the true parameters.
Consider a Gaussian model given by: $Y=w^{*}X+b^{*}+\epsilon$, with $X\sim\mathcal{N}(\mu_{X},\sigma_{X}^{2})$, $\epsilon\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$ independent of $X$, Building on prior work in information theory and statistical learning (e.g.,(Leung and Barron, 2006; Gabrié et al., 2018)), mutual information between input $X$ and prediction $\hat{Y}$, parameterized by model parameter $w$, is expressed as:
(15) $I_{w}(X;\hat{Y})=\frac{1}{2}\ln\left(1+\frac{w\sigma_{X}^{2}}{\sigma_{\epsilon}^{2}}\right).$
[…]